OpenStudy (anonymous):

Proof:

OpenStudy (anonymous):

\[\frac{ \cos³ \alpha +\sin³ \alpha }{ \cos \alpha + \sin \alpha } = 1 - \frac{ 1 }{ 2 } \sin2 \alpha\]

OpenStudy (anonymous):

I was thinking of starting off with the formula of 1/2sin2a = sin a * cos a for the RHS

OpenStudy (anonymous):

but not sure how to get rid of the ³'s in LHS

OpenStudy (yttrium):

\[\frac{ cosx+sinx(\cos^2x-sinxcosx+\sin^x) }{ cosx+sinx }\] \[\cos^2-sinxcosx+\sin^2x\] We know that sin^2x+cos^x = 1 Hence\[1-cosxsinx\] Therefore, \[1-\frac{ 1 }{ 2 }sinxcosx\]

OpenStudy (anonymous):

yu gave complete solution @Yttrium ...not the way of approach only

OpenStudy (anonymous):

No problem, I'm not just copying it. Been working on exercises like these for 2 hours now, got a test this week and I just read it through to understand it.

OpenStudy (yttrium):

I believe @Mieeeel can understand it.

OpenStudy (yttrium):

I mean, easily he can understand. Because he is in there, so far.

OpenStudy (anonymous):

Yea I understand it, sometimes I just lack insight in exercises like these, hence I try making them

OpenStudy (yttrium):

So how's it? Just be honest if you don't understand something.

OpenStudy (anonymous):

I just didn't know how to get started, I was looking for ways to scratch the ³ by the division, which is not allowed because there are no multiplications in the fracture.

OpenStudy (yttrium):

Just review your algebra. :))

OpenStudy (anonymous):

ohh..i just thought it is square term...just noticed that its cubic ...then yu can use the A^3+B^3 identity...applying that and solving progessively ,yu will get RHS

OpenStudy (anonymous):

Wait, I don't get the last step. You get 1 - sinx cos x from the baseformula, I get that But then you go on to 1 - 1/2 (sinx cos x). Where does the 1/2 come from?

OpenStudy (yttrium):

Ohh. It's an error. It should be 1-1/2 (sin2x) It is becase sin2x = 2sinxcosx What we have is just sinxcosx, hence it is 1/2 (sin2x)